OpenStudy (anonymous):

An exponential growth equation is given to you. y=500 × 1.08x. What is the growth rate? How do you know?

OpenStudy (anonymous):

i dont see any exponent to the given equation, is the problem y=500*1.08^x?

OpenStudy (ranga):

The general exponential growth equation is: y = a * (1 + r)^x where a is the initial value and r is the growth rate. Here, y = 500 * 1.08^x = 500(1 + 0.08)^x r = 0.08 The 0.08 is the growth rate. It can also be written as 8% growth rate.

OpenStudy (anonymous):

Yes x is the exponent thing

OpenStudy (the_fizicx99):

The growth rate is 1.08^x where x is any number, y =P(1+r)^x, where P is the principal amount and r is the rate of change in decimal form.

OpenStudy (the_fizicx99):

What ever the rate is you add 1 to it >.<